OpenStudy (anonymous):

Solve for x: one over five (5x + 12) = 18 seventy-eight over five one hundred-two over five 6 30

OpenStudy (luffingsails):

What is the "one over five..." stuff?

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

its a fraction

OpenStudy (igreen):

\(\dfrac{1}{5}(5x + 12) = 18\) We can divide 1/5 to both sides or multiply 5, so what's 18 * 5?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

18*5=90

OpenStudy (igreen):

Yep, so we have: \(5x + 12= 90\) Now subtract 12 to both sides.

OpenStudy (anonymous):

its 78/5

OpenStudy (igreen):

Yep

OpenStudy (igreen):

Divide that

OpenStudy (anonymous):

thankyou

OpenStudy (igreen):

No problem, Happy to Help \(\Huge\ddot\smile\)

OpenStudy (anonymous):

15.6

OpenStudy (igreen):

Yep, that's our answer.